Just-in-Time Inventory The Field Manufacturing Company uses the perpetual inventory system for its raw material inventory. During 2012, Field plans to include raw material costing $2,400,000 in the products that it manufactures. Henry Field, president of the company, wants to adopt the just-in-time manufacturing philosophy for the raw materials inventory during 2012. He wants to have only the raw material needed for the next day's production at the end of each day. The factory operates 300 days each year. Historically, the raw materials inventory balance at the end of the day has averaged $60,000 cost. Field has an annual inventory carrying cost equal to 22% of total inventory cost. Round to the nearest dollar.

Required

a. What is the anticipated inventory carrying cost (in dollars) if Field does not adopt the just-in-time manufacturing philosophy?

b. Calculate the average level (in dollars) for the raw materials inventory if Field adopts the just-in-time manufacturing philosophy.

c. Calculate the reductions in the raw materials inventory level and the raw materials inventory annual carrying cost if Field adopts the just-in-time manufacturing philosophy.

Just-in-time Inventory
a. $
b. $ per day
c. Reduction in materials inventory carrying cost $
Answer #1

ANSWER

a. Anticipated inventory carrying cost= 22% of total inventory cost i.e on 24,00,000

=5,28,000/-

b. If field adopts just-in time inventory average level will be 24,00000/300 days= 8000

c. reduction in raw material level from 60000 t0 8000.

Just in time carrying cost=8000*40=3,20,000/-

perpetual inventory carrying cost=5,28,000/- (As above)

difference=2,08,000/-